OpenStudy (anonymous):

Find the zeros of the function: y= 2x^2 - 3x - 3

OpenStudy (ddcamp):

When \[ax^2 +bx + c = 0\] You can use the quadratic formula to solve for x \[\frac{ -b \pm \sqrt{b^2 - 4ac} }{ 2a }\]

OpenStudy (anonymous):

Yes, but I don't see how that helps me when it comes to finding the zeros(x-intercepts) of that function.

OpenStudy (ddcamp):

Sorry, let me rephrase that: y = ax² + bx + c, with a = 2, b = -3, and c = -3, right?

OpenStudy (anonymous):

Correct.

OpenStudy (ddcamp):

So when y is 0, 0 = ax² + bx + c.

OpenStudy (anonymous):

Oh I see what you mean. Sorry was thinking about this the wrong way.

OpenStudy (ddcamp):

Plugging the a, b, and c values into the quadratic formula will tell you the x values where y is 0.

OpenStudy (anonymous):

Thank you very much for your time